# The $\ell$-test: leveraging sparsity in the Gaussian linear model for improved inference

### v1

@@ Appendix D Characterization of $\hat{\bm{\beta}}^{\lambda}_{-j}(b)$ as a function of $b$
7       Find $d_{2}:=\min\left\{d>0:\textrm{ at least one coordinate of }\hat{\bm{\beta}}^{\lambda}_{S(b)}+d\bm{\gamma}_{S(b)}\textrm{ is }0\right\}$.
@@ 16       Set $b=b+d$
Algorithm 1 Piecewise linear characterization of $\hat{\bm{\beta}}^{\lambda}(b)$ (as defined in (D.1))
In Sections 3 and 4, we saw that for obtaining the $\ell$-test confidence intervals for $\beta_{j}$, one needs to evaluate the function $\hat{\bm{\beta}}^{\lambda}_{-j}(b)$ for different values of $b$. Recall from (2.3) that $\hat{\bm{\beta}}^{\lambda}_{-j}(b)$ is defined as
@@
$\hat{\bm{\beta}}^{\lambda}(b):=\underset{\bm{\beta}}{\arg\min}\left(\frac{1}{2n}\|\bm{y}-b\bm{v}-\bm{Z}\bm{\beta}\|^{2}+\lambda\|\bm{\beta}\|_{1}\right).$ (D.1)
@@
In the following proposition, we first show that $\hat{\bm{\beta}}^{\lambda}(b)$ is a piecewise linear function of $b$, in which we take heavy inspiration from Rosset and Zhu (2007) where the authors establish that the optimizers of $\ell_{1}$-penalized, twice-differentiable likelihoods are piecewise linear in the regularization parameter $\lambda$ and provide algorithm for generating these paths.
Proposition D.1.
Assume that $\bm{y},\bm{v}\in\mathbb{R}^{n}$ and $\bm{Z}\in\mathbb{R}^{n\times d}$, for $n>d$. Then, for $\hat{\bm{\beta}}^{\lambda}(b)$ defined as in (D.1), we have that $\hat{\bm{\beta}}^{\lambda}(b)$ is a continuous, piecewise linear function of $b$.
@@
First note that due to Berge’s Maximum Theorem (Berge, 1963), $b\mapsto\hat{\bm{\beta}}^{\lambda}(b)$ is a continuous transformation. Now for any $b\in\mathbb{R}$, define,
$S(b)=\left\{j\in[1:d]:\left(\hat{\bm{\beta}}^{\lambda}(b)\right)_{j}\neq 0\right\}.$ (D.2)
@@
$\frac{\partial}{\partial b}\hat{\bm{\beta}}_{S(b)}^{\lambda}(b)=-\left(\bm{Z}_{S(b)}^{T}\bm{Z}_{S(b)}\right)^{-1}\bm{Z}_{S(b)}^{T}\bm{v}$ (D.3)
@@
We now turn our attention to methods of exactly generating these piecewise linear paths, following an approach analogous to Rosset and Zhu (2007). To begin with, denote,
@@
[…]
@@
$\hat{\beta}^{\lambda}_{i}(b)\neq 0\Leftrightarrow\left\lvert\left(\partial_{\bm{\beta}}L\big{|}_{\bm{\beta}=\hat{\bm{\beta}}^{\lambda}(b)}\right)_{i}\right\rvert=\lambda\Leftrightarrow\left\lvert\bm{Z}_{i}^{T}\left(\bm{y}-\bm{v}b-\bm{Z}\hat{\bm{\beta}}^{\lambda}(b)\right)\right\rvert=n\lambda.$ (D.4)
We use Equation (D.3), (D.4), along with Proposition D.1 and the fact that the set $S(b)$ (from (D.2)) changes only at non-differentiability points of $\hat{\bm{\beta}}(b)$ to devise an algorithm to exactly generate the paths of $\hat{\bm{\beta}}(b)$ as a function of $b$ in Algorithm 1. The algorithm uses a notation defined below for the ‘derivative’ of $\hat{\bm{\beta}}^{\lambda}(b)$:
@@
Appendix E Choice of the tuning parameter, $\lambda$
E.1 The min rule vs. The 1se rule
In Section 2.4, we justified that a reasonable way of choosing $\lambda$ for testing $H_{j}$ so that it does not invalidate our theory surrounding the $\ell$-test can be to cross validate on $(\tilde{\bm{y}},\bm{X}_{-j})$, where $\tilde{\bm{y}}$ is drawn from the conditional distribution of $\bm{y}\mid\bm{S}^{(j)}$, under $H_{j}$. For computational convenience, we have used cross-validation on $(\bm{X}_{-j},\bm{\tilde{y}})$ instead of $(\bm{X},\bm{\tilde{y}})$, as the two choices result in almost identical performances of the resulting methods and because the LASSO estimates obtained using $(\bm{X}_{-j},\bm{y})$ is the same as that using $(\bm{X}_{-j},\bm{\tilde{y}})$, this enables us to recycle some common information from the latter dataset, thereby saving computation time while obtaining the conditional distribution of $\hat{\beta}_{j}\mid\bm{S}^{(j)}$, under $H_{j}$.
@@ Appendix E Choice of the tuning parameter, $\lambda$
In Section F.3, we provide the empirical coverage and lengths of the resulting confidence intervals when using the 1se rule for selecting $\lambda$. To summarize our findings, we observe that the 1se rule and the min rule have similar performances for constructing $\ell$-test confidence intervals except for the case when the $\bm{\beta}$ is very dense, in which case the min rule results in intervals of shorter length. Hence, we recommend using the min rule for choosing $\lambda$.
E.2 The randomness in the choice of $\lambda$
The rule for the choice of $\lambda$ we discussed involves sampling, $\tilde{\bm{y}}\sim\bm{y}\mid\bm{S}^{(j)}$, under $H_{j}$ and then running cross-validation on $(\bm{X}_{-j},\tilde{\bm{y}})$. This suggests towards some inherent sources of variability in the method—the random sampling of $\tilde{\bm{y}}$ and the random splitting of the dataset to perform cross-validation. In order to understand its effect, we perform $m=100$ replications of an experiment under the setting of the left panel of Figure 2 where for each replicate we form a linear model with this design matrix and obtain p-value for testing $H_{j}:\beta_{j}=0$ for $m=100$ samples of $\tilde{\bm{y}}$. Let $p_{ij}$ denote the p-value after sampling $\tilde{\bm{y}}$ for the $j^{\mathrm{th}}$ time for the dataset $(\bm{y}^{(i)},\bm{X}^{(i)})$ (corresponding to the $i^{\mathrm{th}}$ replication of drawing $(\bm{y},\bm{X})$ from a pre-determined distribution). For these $m^{2}=10^{4}$ p-values, we plot the empirical estimate of the overall standard deviation of the p-values, $\sqrt{\mathrm{Var}(p_{ij})}$ against the standard deviation conditioned on a replicate, $\sqrt{\mathbb{E}\mathrm{Var}\left(p_{ij}\mid(\bm{y}^{(i)},\bm{X}^{(i)})\right)}$ in Figure 6 in the log-scale. These quantities are estimated using $\sqrt{\frac{1}{m^{2}-1}\left(\sum_{i,j=1}^{m}(p_{ij}-\bar{p}_{\cdot\cdot})^{2}\right)}$ and $\sqrt{\frac{1}{m^{2}-m}\left(\sum_{i,j=1}^{m}(p_{ij}-\bar{p}_{i\cdot})^{2}\right)}$, respectively, where, $\bar{p}_{i\cdot}$ represents the mean of the entries in $\{p_{ij}:1\leq j\leq m\}$, while $\bar{p}_{\cdot\cdot}$ represents the mean of all the p-values, $\{p_{ij}:1\leq i,j\leq m\}$.
@@ E.2 The randomness in the choice of $\lambda$
Appendix F Further Experiments
F.1 Performance of the $\ell$-test under different settings
@@
F.2 Robustness of the $\ell$-test to violations of the linear model assumptions
In this section, we extend the results in Section 5.2 by performing more extensive experiments to empirically evaluate the robustness of the validity of the $\ell$-test to the violations in the assumptions of the Gaussian linear model. We will be under the same exact setup of Section 5.2 and will consider the following settings, each aimed at testing a specific kind of violation.
  • Setting 1 (Violation of the Gaussianity of errors—effect of heavy tails): For each specific value of $(n,d)$, we draw i.i.d. errors from a $t$ distribution with $\nu$ degrees of freedom. We vary $\nu$ between 30 and 2. For $\nu>2$, we also standardize the mean zero errors with the standard deviation of the $t_{\nu}$ distribution. We do not do this for $t_{2}$ as it does not have a finite second moment. As $\nu$ varies from 30 to 2, the tails of $t_{\nu}$ gets fatter as compared to the normal distribution. We summarize our results in Figure 9.
  • Setting 2 (Violation of the Gaussianity of errors—effect of skewness): We consider a setup similar to that in Setting 1 but instead consider Gamma distributed error with scale parameter 1 and shape parameter, $\alpha$. We vary $\alpha$ between 1 and 10 and for each error draw, we standardize the error with the mean and standard deviation of the Gamma$(1,\alpha)$ distribution. This error distribution is asymmetric for smaller values of $\alpha$ and moves towards symmetry as the value of $\alpha$ increases. We summarize our results in Figure 10.
  • Setting 3 (Violation of homoskedasticity of error): We again consider a similar setup as above, but change the error distribution as follows: For design matrix, $\bm{X}$, we define $m_{\bm{X}}$ to be the median of the mean of the rows, that is median of the elements of $\bm{X}\bm{1}/d$. Let $r_{i}$ denote mean of the $i^{\mathrm{th}}$ row of the design matrix, we generate the error vector $\bm{\epsilon}=(\epsilon_{1},\dots,\epsilon_{n})^{T}$, where,
    $\epsilon_{i}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}\begin{cases}\mathcal{N}(0,1),&\textrm{ if }r_{i}\leq m_{\bm{X}}\\ \mathcal{N}(0,\eta^{2}),&\textrm{ if }r_{i}>m_{\bm{X}}\end{cases},$
    where $\eta^{2}>0$ is a quantity, specified by us, that controls the heteroskedasticity in the error term. For $\eta=1$, the distribution is homoskedastic while becomes heteroskedastic for larger and smaller values of $\eta$. We vary $\eta^{2}$ in the set $\{0.01,0.25,0.5,1,4,8\}$ and compare the performance of the two tests for each of these values. The results are summarized in Figure 11.
@@ F.2 Robustness of the $\ell$-test to violations of the linear model assumptions
    where $\delta$ is variable we will control, and $\bm{X}_{i}^{\delta}$ denotes a vector whose $j^{\mathrm{th}}$ entry is the $j^{\mathrm{th}}$ entry of $\bm{X}_{i}$ raised to the exponent, $\delta$. $\delta=1$ recovers the usual linear model and larger departure of $\delta$ from 1 imparts higher degree of non-linearity to the model. We vary $\delta$ in the set, $\{0.3,0.5,1,2,3,4\}$ and summarize the results in Figure 12.
The results from all the three simulations suggest that the $\ell$-test and the $t$-test have similar degree of tolerance against violations of the linear model assumptions. The $t$-test exhibits robustness against the violation of the Gaussianity assumption in the error term, as Figures 9 and 10 indicate and we see a similar behavior for the $\ell$-test. Notably, the performance of $\ell$-test is almost similar to that of the $t$-test in the extreme cases, such as when the degrees of freedom for the error $t$-distribution is 2 in Figure 9 (indicating fat tails of the error distribution) or when the shape parameter of the centered gamma distributed error is 1 (which essentially is a centered exponential distribution with rate 1, indicating a high degree of skeweness in the error term). For Setting 3 with heteroskedastic errors, we see from Figure 11 that the $\ell$-test’s size does depart from its nominal target of $0.05$ as $\eta$ moves away from 1 (which is the homoskedastic case). However, this departure is of a similar degree as that of the $t$-test, so that both the tests exhibit similar robustness properties under this setting. From Figure 12 as well, we see that both the $t$-test and the $\ell$-test are robust to the violation of non-linearity.
Figure 9: Effect of $t$-distributed errors on the size of the $\ell$-test and the $t$-test, for different degrees of freedom. The error bars represent plus or minus two standard errors.
Figure 10: Effect of Gamma distributed errors (with scale parameter 1) on the size of the $\ell$-test and the $t$-test, for different values of the shape parameter. The error bars represent plus or minus two standard errors.
Figure 11: Effect of heteroskedasticity on the size of the $\ell$-test and the $t$-test. Values of $\eta$ away from 1 indicate higher degree of heteroskedasticity. The error bars represent plus or minus two standard errors.
Figure 12: Effect of non-linearity on the size of the $\ell$- and the $t$-test. Departure of the exponent, $\delta$ away from 1 indicates higher degree of non-linearity in the model. The error bars represent plus or minus two standard errors.
F.3 Comparison of the various confidence intervals in the linear model
@@
Figure 13: Length and coverage of various $95\%$ confidence intervals. We set $n=100,d=50,k=5,\bm{\Sigma}=\bm{I},\sigma=1$ and vary $A$. The error bars represent plus or minus two standard errors.
Figure 14: Length and coverage of various $95\%$ confidence intervals. We set $n=100,d=90,k=5,\bm{\Sigma}=\bm{I},\sigma=1$ and vary $A$. The error bars represent plus or minus two standard errors.
@@
F.4 LASSO selection-adjusted confidence sets for linear model coefficients
Expanding on Section 5.4, we compare the various conditional-on-LASSO-selection inference methods under another setting where $d=90$, with the results plotted in Figure 17. As in Figure 5, there is practically no difference between the performance of $\underline{\hat{C}}_{j}^{\lambda}$ and $\underline{\hat{C}}_{j}^{*\lambda}$ but, unlike in Figure 5, Liu et al. (2018) performs quite similarly to them as well. The reason for this is that, as noted in Section 1.3, Liu et al. (2018)’s method assumes known $\sigma^{2}$ (and hence we provide it the true value of $\sigma^{2}$ in both this simulation and that of Figure 5) while the $\ell$-test does not. Unknown $\sigma^{2}$ is what necessitates the second component of the sufficient statistic $\bm{S}^{(j)}$, namely, $\bm{y}^{T}\bm{y}$, which the $\ell$-test conditions on but Liu et al. (2018)’s method does not. This difference becomes information-theoretically more pronounced as $d$ approaches $n$, as it does in the simulation in Figure 17, because the residual degrees of freedom decreases, resulting in a relative power loss for the $\ell$-test-based confidence intervals that approximately offsets the benefit of the $\ell$-test. To confirm this explanation, we can easily derive a version of the $\ell$-test and corresponding conditional confidence intervals that assumes $\sigma^{2}$ is known and hence does not condition on $\bm{y}^{T}\bm{y}$ (see footnote in Section 6). This is plotted as dashed curves in Figure 17, and as expected these methods again provide consistently shorter confidence intervals than the method in Liu et al. (2018).

### v2

@@ Appendix F Characterization of $\hat{\bm{\beta}}^{\lambda}_{-j}(b)$ as a function of $b$
[…]
@@ 16  Set $b=b+d$
Algorithm 2 Piecewise linear characterization of $\hat{\bm{\beta}}^{\lambda}(b)$ (as defined in (F.1))
In Sections 4 and 5, we saw that for obtaining the $\ell$-test confidence intervals for $\beta_{j}$, one needs to evaluate the function $\hat{\bm{\beta}}^{\lambda}_{-j}(b)$ for different values of $b$. Recall from (2.3) that $\hat{\bm{\beta}}^{\lambda}_{-j}(b)$ is defined as
@@
$\hat{\bm{\beta}}^{\lambda}(b):=\underset{\bm{\beta}}{\arg\min}\left(\frac{1}{2n}\|\bm{y}-b\bm{v}-\bm{Z}\bm{\beta}\|^{2}+\lambda\|\bm{\beta}\|_{1}\right).$ (F.1)
@@
In the following proposition, we first show that $\hat{\bm{\beta}}^{\lambda}(b)$ is a piecewise linear function of $b$, in which we take heavy inspiration from piecewiselinear where the authors establish that the optimizers of $\ell_{1}$-penalized, twice-differentiable likelihoods are piecewise linear in the regularization parameter $\lambda$ and provide algorithm for generating these paths.
Proposition F.1.
Assume that $\bm{y},\bm{v}\in\mathbb{R}^{n}$ and $\bm{Z}\in\mathbb{R}^{n\times d}$, for $n>d$. Then, for $\hat{\bm{\beta}}^{\lambda}(b)$ defined as in (F.1), we have that $\hat{\bm{\beta}}^{\lambda}(b)$ is a continuous, piecewise linear function of $b$.
@@
First note that we can use arguments similar to the ones presented in Lemma 9 to conclude that $b\mapsto\hat{\bm{\beta}}^{\lambda}(b)$ is a continuous transformation. Now for any $b\in\mathbb{R}$, define,
$S(b)=\left\{j\in[1:d]:\left(\hat{\bm{\beta}}^{\lambda}(b)\right)_{j}\neq 0\right\}.$ (F.2)
@@
$\frac{\partial}{\partial b}\hat{\bm{\beta}}_{S(b)}^{\lambda}(b)=-\left(\bm{Z}_{S(b)}^{T}\bm{Z}_{S(b)}\right)^{-1}\bm{Z}_{S(b)}^{T}\bm{v}$ (F.3)
@@
We now turn our attention to methods of exactly generating these piecewise linear paths, following an approach analogous to piecewiselinear. To begin with, denote,
@@
$\textrm{Minimize: }L(\bm{\beta}^{+}-\bm{\beta}^{-})+\lambda\sum(\beta_{i}^{+}-\beta_{i}^{-})$
$\textrm{Subject to: }\beta_{i}^{+},\beta_{i}^{-}\geq 0,\forall i.$
@@
$\hat{\beta}^{\lambda}_{i}(b)\neq 0\Leftrightarrow\left\lvert\left(\partial_{\bm{\beta}}L\big|_{\bm{\beta}=\hat{\bm{\beta}}^{\lambda}(b)}\right)_{i}\right\rvert=\lambda\Leftrightarrow\left\lvert\bm{Z}_{i}^{T}\left(\bm{y}-\bm{v}b-\bm{Z}\hat{\bm{\beta}}^{\lambda}(b)\right)\right\rvert=n\lambda.$ (F.4)
We use Equation (F.3), (F.4), along with Proposition F.1 and the fact that the set $S(b)$ (from (F.2)) changes only at non-differentiability points of $\hat{\bm{\beta}}(b)$ to devise an algorithm to exactly generate the paths of $\hat{\bm{\beta}}(b)$ as a function of $b$ in Algorithm 2. The algorithm uses a notation defined below for the ‘derivative’ of $\hat{\bm{\beta}}^{\lambda}(b)$:
@@
$:=$ $\bm{0},\textrm{ otherwise}.$
Appendix G Choice of the tuning parameter, $\lambda$
G.1 The min rule vs. The 1se rule
@@ Appendix G Choice of the tuning parameter, $\lambda$
In Section H.3, we provide the empirical coverage and lengths of the resulting confidence intervals when using the 1se rule for selecting $\lambda$. To summarize our findings, we observe that the 1se rule and the min rule have similar performances for constructing $\ell$-test confidence intervals except for the case when the $\bm{\beta}$ is very dense, in which case the min rule results in intervals of shorter length. Hence, we recommend using the min rule for choosing $\lambda$.
G.2 The randomness in the choice of $\lambda$
[…]
@@ G.2 The randomness in the choice of $\lambda$
Appendix H Further Experiments
H.1 Performance of the $\ell$-test under different settings
@@
H.2 Robustness of the $\ell$-test to violations of the linear model assumptions
In this section, we extend the results in Section 6.2 by performing more extensive experiments to empirically evaluate the robustness of the validity of the $\ell$-test to the violations in the assumptions of the Gaussian linear model. Aside from the details of the model violations described below, the setup is identical to that of Section 6.2.
  • Setting 1 (Violation of the Gaussianity of errors—effect of heavy tails): For each specific value of $(n,d)$, we draw i.i.d. errors from a $t$-distribution with $\nu$ degrees of freedom. We vary $\nu$ between 30 and 2. For $\nu>2$, we also standardize the mean zero errors with the standard deviation of the $t_{\nu}$ distribution. We do not do this for $t_{2}$ as it does not have a finite second moment. As $\nu$ varies from 30 to 2, the tails of $t_{\nu}$ gets fatter as compared to the normal distribution. We summarize our results in Figure 13 (corresponding to $\rho=0$) and Figure 17 (corresponding to $\rho=0.5$).
[…]
  • Setting 3 (Violation of homoskedasticity of error): We again consider a similar setup as above, but change the error distribution as follows: For design matrix, $\bm{X}$, let $r_{i}$ denote mean of the $i^{\mathrm{th}}$ row of the design matrix. We generate the error vector $\bm{\epsilon}=(\epsilon_{1},\dots,\epsilon_{n})^{T}$, based on the following rule:
    $\epsilon_{i}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}\begin{cases}\mathcal{N}(0,1),&\textrm{ if }r_{i}\leq 0\\ \mathcal{N}(0,\eta^{2}),&\textrm{ if }r_{i}>0\end{cases}.$ (H.1)
    where $\eta^{2}>0$ is a quantity, specified by us, that controls the heteroskedasticity in the error term. The scheme (H.1) represents a simple way of imparting heteroskedasticity, where the error variance of a row depends on the mean value of its entries. For $\eta=1$, the distribution is homoskedastic while becomes heteroskedastic for larger and smaller values of $\eta$. We vary $\eta^{2}$ in the set $\{0.01,0.25,0.5,1,4,8\}$ and compare the performance of the two tests for each of these values. The results are summarized in Figure 15 (corresponding to $\rho=0$) and Figure 19 (corresponding to $\rho=0.5$).
@@ H.2 Robustness of the $\ell$-test to violations of the linear model assumptions
    where $\delta$ is variable we will control, and $\bm{X}_{i}^{\delta}$ denotes a vector whose $j^{\mathrm{th}}$ entry is the absolute value of the $j^{\mathrm{th}}$ entry of $\bm{X}_{i}$ raised to the exponent, $\delta$, and then multiplied back by its sign. $\delta=1$ recovers the usual linear model and larger departure of $\delta$ from 1 imparts higher degree of non-linearity to the model. We vary $\delta$ in the set, $\{0.3,0.5,1,2,3,4\}$ and summarize the results in Figure 16 (corresponding to $\rho=0$) and in Figure 20 (corresponding to $\rho=0.5$).
The results from all the four simulations suggest that the $\ell$-test and the $t$-test have similar degree of tolerance against violations of the linear model assumptions. The $t$-test exhibits robustness against the violation of the Gaussianity assumption in the error term, as Figures 13 and 14 indicate and we see a similar behavior for the $\ell$-test. Notably, the performance of $\ell$-test is almost similar to that of the $t$-test in the extreme cases, such as when the degrees of freedom for the error $t$-distribution is 2 in Figure 13 (indicating fat tails of the error distribution) or when the shape parameter of the centerd gamma distributed error is 1 (which essentially is a centerd exponential distribution with rate 1, indicating a high degree of skeweness in the error term). For Setting 3 with heteroskedastic errors, we see from Figure 15 that the $\ell$-test’s size does depart from its nominal target of $0.05$ as $\eta$ moves away from 1 (which is the homoskedastic case). However, this departure is of a similar degree as that of the $t$-test, so that both the tests exhibit similar robustness properties under this setting. From Figure 16 as well, we see that both the $t$-test and the $\ell$-test are robust to the violation of non-linearity. Furthermore, Figures 17, 18, 19 and 20 show that the conclusions remain valid even when inter-variable dependencies are introduced with $\rho=0.5$.
Figure 13: $\rho=0$. Effect of $t$-distributed errors on the size of the $\ell$-test and the $t$-test, for different degrees of freedom. The error bars represent plus or minus two standard errors.
Figure 14: $\rho=0$. Effect of Gamma distributed errors (with scale parameter 1) on the size of the $\ell$-test and the $t$-test, for different values of the shape parameter. The error bars represent plus or minus two standard errors.
Figure 15: $\rho=0$. Effect of heteroskedasticity on the size of the $\ell$-test and the $t$-test. Values of $\eta$ away from 1 indicate higher degree of heteroskedasticity. The error bars represent plus or minus two standard errors.
Figure 16: $\rho=0$. Effect of non-linearity on the size of the $\ell$- and the $t$-test. Departure of the exponent, $\delta$ away from 1 indicates higher degree of non-linearity in the model. The error bars represent plus or minus two standard errors.
Figure 17: Same as Figure 13 but with $\rho=0.5$.
Figure 18: Same as Figure 14 but with $\rho=0.5$.
Figure 19: Same as Figure 15 but with $\rho=0.5$.
Figure 20: Same as Figure 16 but with $\rho=0.5$.
H.3 Comparison of the various confidence intervals in the linear model
@@
Figure 21: Length and coverage of various $95\%$ confidence intervals. We set $n=100,d=50,k=5,\bm{\Sigma}=\bm{I},\sigma=1$ and vary AA. The error bars represent plus or minus two standard errors.
Figure 22: Length and coverage of various $95\%$ confidence intervals. We set $n=100,d=90,k=5,\bm{\Sigma}=\bm{I},\sigma=1$ and vary AA. The error bars represent plus or minus two standard errors.
@@
H.4 LASSO selection-adjusted confidence sets for linear model coefficients
Expanding on Section 6.4, we compare the various conditional-on-LASSO-selection inference methods under another setting where $d=90$, with the results plotted in Figure 25. As in Figure 5, there is practically no difference between the performance of $\underline{\hat{C}}_{j}^{\lambda}$ and $\underline{\hat{C}}_{j}^{*\lambda}$ but, unlike in Figure 5, liu2018powerful performs quite similarly to them as well. The reason for this is that, as noted in Section 1.3, liu2018powerful’s method assumes known $\sigma^{2}$ (and hence we provide it the true value of $\sigma^{2}$ in both this simulation and that of Figure 5) while the $\ell$-test does not. Unknown $\sigma^{2}$ is what necessitates the second component of the sufficient statistic $\bm{S}^{(j)}$, namely, $\bm{y}^{T}\bm{y}$, which the $\ell$-test conditions on but liu2018powerful’s method does not. This difference becomes information-theoretically more pronounced as $d$ approaches $n$, as it does in the simulation in Figure 25, because the residual degrees of freedom decreases, resulting in a relative power loss for the $\ell$-test-based confidence intervals that approximately offsets the benefit of the $\ell$-test. To confirm this explanation, we can easily derive a version of the $\ell$-test and corresponding conditional confidence intervals that assumes $\sigma^{2}$ is known and hence does not condition on $\bm{y}^{T}\bm{y}$ (see footnote in Section 7). This is plotted as dashed curves in Figure 25, and as expected these methods again provide consistently shorter confidence intervals than the method in liu2018powerful.